OpenStudy (anonymous):

integrate 3+5t dt, find the C also

OpenStudy (anonymous):

The gas piston of an M4 carbine rifle with in initial position of x0=5[mm].experiences a velocity defined by x’=3+5t for one second. What is the piston’s finale position?

OpenStudy (anonymous):

Are we using the definition of an integral, or are we just taking an anti derivative?

OpenStudy (anonymous):

Okay so do you know how to find the indefinite integral here? \[\int 3+5t\ dt\]

OpenStudy (anonymous):

of course it will be 3t+5t square/2 + c, not the point is how to find the C

OpenStudy (anonymous):

did you ready the question plz help me.

OpenStudy (anonymous):

The gas piston of an M4 carbine rifle with in initial position of x0=5[mm].experiences a velocity defined by x’=3+5t for one second. What is the piston’s finale position?

OpenStudy (anonymous):

Okay, so we know that \[5 = 3(0)+5(0)^2+c\]Can you solve for c?

OpenStudy (anonymous):

C will be 5

OpenStudy (anonymous):

So you have the function of position with respect to time. They gave you a time.

OpenStudy (anonymous):

So what is the final position, please write the full step

OpenStudy (anonymous):

the time is in one secant

OpenStudy (anonymous):

\[3(1)+5(1)^2+5\]Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

My Dear thanks for helping me please once again ready the question and write, the step and final anwer

OpenStudy (anonymous):

I mean read the question sorry by mistake it wrote ready

OpenStudy (anonymous):

you mean the final answer is 13 mm

OpenStudy (anonymous):

Yeah....

OpenStudy (anonymous):

I am studying dynamics of structure that is why

OpenStudy (anonymous):

how you find 5=3(0)+5(0)2+c, zero in here where zero come from 0

OpenStudy (anonymous):

We let \(t=0\) at the initial time. And \(t=1\) at the final time.

OpenStudy (anonymous):

Hey wio thanks alot i got the answer but i need the steps

OpenStudy (anonymous):

please write from start to end

OpenStudy (anonymous):

To find c, you need a sample point. They point they gave you is (0,5).

OpenStudy (anonymous):

" initial position of x0=5[mm]" means (0,5)

OpenStudy (anonymous):

then

OpenStudy (anonymous):

suppose someone gave you this question

OpenStudy (anonymous):

The gas piston of an M4 carbine rifle with in initial position of x0=5[mm].experiences a velocity defined by x’=3+5t for one second. What is the piston’s finale position?

OpenStudy (anonymous):

from where you start the solution

OpenStudy (anonymous):

hey anyone plz help me

OpenStudy (anonymous):

see we integrate 3+5t dt = 3t+5t square / 2 + c

OpenStudy (anonymous):

how to find C